A tank contains 440 litres of a mixture which is 23.0% acid. How many litres of a mixture which is 12.0% acid must be added to get a mixture which is 14.5% acid.
Found 2 solutions by checkley77, nerdybill:
Answer by checkley77(12844)

.23*440+.12x=.145(440+x)
101.2+.12x=63.8+.145x
.12x-.145x=63.8-101.2
-.025x=-37.4
x=-37.4/-.025
x=1,496 liters of 12% acid are used.
Proof:
.23*440+.12*1,496=.145(440+1,496)
101.2+179.52=.145*1,936
280.72=280.72

Answer by nerdybill(7384)

Let x = amount of 12% acid added
.
"amt of acid from 23%" + "amt of acid from 12%" = "amt in final mixture"
.23(440) + .12x = .145(440+x)
101.2 + .12x = 63.8 + .145x
37.4 + .12x = .145x
37.4 = 0.025x
1496 liters = x